Consider two industries, industry W and industry X. In industry W there are five companies, each with a market share of 20% of total sales. In industry X, there are six companies. One company has a 50% market share and each of the other five firms has a market share of 10%.

a.   Calculate the four-firm concentration ratio for each industry.

b.   Calculate the Herfindahl-Hirschman Index (HHI) for each industry.

c.   What do the values of the two concentration measures imply about the degree of market power in the two industries?
Answer #1

A) concentration ratio=share of first 4 leadings firm /total market

Industry W first four leading firms have share equals to 4*20=80 and total market share is equal to 20*5=100

Industry w first four leading firma have share equals to 80(50+10+10+10), and total market share is 100(50+10+10+10+10+10)

industry W concentration ratio=80/100=0.8

Industry X CONCENTRATION RATION =80/100=0.8

B)HHI=10,000(sum of square of each firm's market share)

For Industry W,

Each firms have same market share equal to 0.2

So,

HHI=10,000(5*(0.2)^2)=10,000(5*0.04)=2000

For Industry X,

Leading firm have Market share of 0.5 and rest five firms have 0.1 .

So,

HHI=10,000(0.5^2+5*(0.1)^2)

HHI=10,000*0.30=3000

C) concentration ratio tells In both industry market power is same and high as both industry has same concentration ratio ,0.8.

But HHI index tells market power in industry X is more than industry W as X has 3000 HHI while W has just 2000HHI. Higher the HHI ,higher the market power.